Problem 6-5 Julia Baker died, leaving to her husband Adams an insurance policy contract that provides that the beneficiary (Adams) can choose any one of the following four options. Money is worth 2.50% per quarter, compounded quarterly. Compute Present value if:

A: $4,090 every 3 months payable at the end of each quarter for 5 years.

B: $19,350 immediate cash and $1,935 every 3 months for 10 years, payable at the beginning of each 3-month period.

C: $4,090 every 3 months for 3 years and $1,420 each quarter for the following 25 quarters, all payments payable at the end of each quarter.
Answer #1
Please hit LIKE button if this helped. For any further explanation, please put your query in comment, will get back to you.
Part A:
Interest Rate for Quarter 2.50%
n (5 Years*4 Quarters) 20 Quarters
Amount at end of every quarter $         4,090
PVAF n=1-20, i=2.5%        15.5892
Present Value 4090*15.5892 $      63,760
Part B:
Interest Rate for Quarter 2.50%
n (10 Years*4 Quarters) 40 Quarters
Amount at the beginning of Quarter $         1,935
PVAF n=0-39, i=2.5%        25.7303
Present Value 1935*25.7303 $      49,788
Add: Immediate Payment $      19,350
Total Present Value $      69,138
Part C:
Interest Rate for Quarter 2.50%
n (3 Years*4 Quarters) 12 Quarters
Amount at end of every quarter $         4,090
PVAF n=1-12, i=2.5%        10.2578
Present Value 4090*10.2578 $      41,954 A
n 25 Quarters
Amount at end of every quarter $         1,420
PVAF n=13-37, i=2.5%        13.6996
Present Value 1420*13.6996 $      19,453 B
Total Present Value (A+B) $      61,408
Summary:
Present Value of:
Option A $      63,760
Option B $      69,138 Best
Option C $      61,408
Answer #2

Q2: Rajeev Gandhi died, leaving to his wife an insurance policy contract that provides beneficiary (wife) can choose any one of the following four options.

 a. $55,000 immediate cash.

 b. $4,000 every 3 months payable at the end of each quarter for 5 years.

 c. $18,000 immediate cash and $1,800 every 3 months for 10 years, payable at the beginning of each

3-month period.

 d. $4,000 every 3 months for 3 years and $1,500 each quarter for the following 25 quarters, all payments payable at the end of each quarter.

Instructions

If annual interest rate 10% per, compounded quarterly, which option would you recommend that

the wife exercise?


source: intermediate accounting i
answered by: Anas Masri